OpenStudy (anonymous):

Simplify the expression 5^3 × 5^8.

OpenStudy (anonymous):

(5x5x5)x(5x5x5x5x5x5x5x5)

OpenStudy (anonymous):

=48828125

OpenStudy (acxbox22):

in general \[a^b \times a^c = a^{b+c}\]

OpenStudy (anonymous):

no @kathyzhu23

OpenStudy (acxbox22):

when multiplying powers with the same base add the powers

OpenStudy (vlery):

5^11

OpenStudy (anonymous):

5^11

OpenStudy (anonymous):

thats easy. just add the exponents.

OpenStudy (acxbox22):

so 5^3 x 5^8 = 5^(3+8) =5^11